alebra:quare root of 24
Answer by Theo(13342)

apparently you need to simplify this without the aid of a calculator.

sqrt(24) = sqrt(2*12) = sqrt(2*2*6) = sqrt(2*2*2*3) = 2*sqrt(6).

I don't think it can be simplified any further.

another option would be:

sqrt(24) = sqrt(3*8) = sqrt(3*2*4) = sqrt(3*2*2*2) = 2*sqrt(3*2) = 2*sqrt(6).

we get the same answer.